OpenStudy (anonymous):

sketch the graph of the function and its inverse on the same coordinate plane: y=x^3 + 2

OpenStudy (anonymous):

dude cookies is diggin yo shirt!!

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

lol i meant pic haha it is on a shirt to!!

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

no i luv adventure time soooo very childish of me but hey rainicorns are very rare you know

OpenStudy (anonymous):

for inverse you switch x and y so x = y^3 +2. x = y^3 + 2 x - 2 = y^3 y = (x-2)^1/3 and graph that and the original function by plotting points

OpenStudy (anonymous):

yes... those rainicorns are very hard to find... believe me, I've tried looking